# AgentsCoMerge: Large Language Model Empowered Collaborative Decision Making for Multi-Lane Merging

### v1

@@ 5.2 Implementation Details
We leverage Qwen-7B [59] as our pre-trained language model, which contains 7.7 billion parameters. For closed-loop evaluation, we set the weights of CS, ES, and SS as $w_{1}=0.25$, $w_{2}=0.25$, and $w_{3}=0.5$, respectively. The penalty factors for collisions and speed violations are set on $\alpha=0.6$ and $\beta=0.9$, respectively.
Figure 7: Qualitative analysis and prompts visualization.
@@ 5.3 Comparison Analysis
Open-loop Evaluation. To further demonstrate the effectiveness of our AgentsCoMerge, and its generalization ability to other scenarios, and real-world scenarios, we evaluate it on the nuScenes dataset. We compare our method with the following methods, NMP [51], SA-NMP [51], FF [52], EO [53], ST-P3 [54], UniAD [55], and GPT-Driver [56]. The results are shown in Table I. As we can see, our proposed method outperforms all the other motion planning baselines on the test set of the nuScenes dataset on both L2 error and collision rate. Specifically, AgentsCoMerge achieves an average L2 error of 0.59m, which is 42.71% lower than the best none-LLM-based method UniAD. For LLM-based methods GPT-Driver which is the second-best method, our method achieves 29.76% lower on L2 error. In terms of collision rate, our method achieves an average collision rate of 0.29%. It is worth noting that our method achieves a collision rate of 0.00% in the 1s prediction horizon, which means that in the 1s prediction waypoints, our method can avoid collisions with other objects in the scene. These comparative analysis results demonstrate the effectiveness of our proposed method in the open-loop evaluation. In addition, the results also show the generalization ability and underline the importance of the powerful hierarchical planning ability.

### v2

@@ 5.2 Implementation Details
We leverage Qwen-7B [63] as our pre-trained language model, which contains 7.7 billion parameters. For closed-loop evaluation, we set the weights of CS, ES, and SS as $k_{1}=0.25$, $k_{2}=0.25$, and $k_{3}=0.5$, respectively. The penalty factors for collisions and speed violations are set on $\alpha=0.6$ and $\beta=0.9$, respectively. In addition, we set the TTC threshold in Eq. 35 as $t_{t}=5s$.
@@ 5.3 Comparison Analysis
Figure 7: Qualitative analysis and prompts visualization.
Open-loop Evaluation. To further demonstrate the effectiveness of our AgentsCoMerge, and its generalization ability to other scenarios, and real-world scenarios, we evaluate it on the nuScenes dataset. We compare our method with the following methods, NMP [50], SA-NMP [50], FF [51], EO [52], ST-P3 [53], UniAD [54], and GPT-Driver [48]. The results are shown in Table I. Our method outperforms all baselines on both metrics. AgentsCoMerge achieves an average L2 error of 0.59m, which is 42.71% lower than UniAD (best none-LLM method) and 29.76% lower than GPT-Driver (best LLM method). For collision rate, our method achieves 0.29% on average, with 0.00% in the 1s prediction horizon. These results demonstrate both the effectiveness of our method and its strong generalization ability through hierarchical planning. In addition, we evaluate the performance on HighD dataset [58] with four other methods including CS-LSTM [59], Dual-Trans [60], WSiP [61], and Social GAN [62]. As shown in Table V, the overall average RMSE of 0.73m of our method represents a 21.5% improvement over the next best baseline, validating our approach’s effectiveness for trajectory prediction in real highway scenarios.
@@ 5.5 Case Study
6 Limitations and Future Work
In this section, we discuss the limitations of our proposed framework. We acknowledge that our current implementation using LLMs faces some computational challenges in terms of computational cost and real-time deployment. The LLM models have billions of parameters, leading to high computational resource requirements and an average decision time of around 10 seconds per round. While this current decision latency is not suitable for real-time autonomous driving, this limitation is primarily technical rather than fundamental. Recent advances show promising directions for optimization through model compression and acceleration techniques [64, 65], as well as quantization and tensor operator fusion optimization [66, 67]. These optimizations have successfully reduced decision times from seconds to milliseconds. The practical viability of LLM-based approaches has been demonstrated by some automotive companies (e.g., Li Auto) that have successfully deployed LLM-based decision systems in autonomous vehicles [68]. While computational optimization is important, it is out of the scope of this paper. Our primary contribution is the design of a multi-agent collaborative decision-making framework that improves autonomous vehicles’ ability to make driving and coordinated decisions. The computational aspects can be addressed through ongoing optimization research.